OpenStudy (anonymous):

y = 8 - 2x Can you make an input/output table for the function, please? Use 0, 1, 2, 3, 4, and 5 as values for x. Thank Ya' very much~

OpenStudy (precal):

sub x=0 y=8-2(0) y=8-0 y=8 do the rest like this